OpenStudy (anonymous):

A cubic foot of water weighs 62.4 lb. A basketball has a radius of 0.39 ft. About how much would the basketball weigh if it was filled with water?

OpenStudy (anonymous):

Weight = 62.4*Vol pounds

OpenStudy (anonymous):

do you know equation for the volume of a sphere?

OpenStudy (anonymous):

\[V = 4/3\pi r ^{3}\]